OpenStudy (anonymous):

Simplify √49 ----- √2

jimthompson5910 (jim_thompson5910):

First take the square root of 49 to get 7 \[\Large \frac{\sqrt{49}}{\sqrt{2}}\] \[\Large \frac{7}{\sqrt{2}}\]

jimthompson5910 (jim_thompson5910):

Now multiply both top and bottom by \(\large \sqrt{2}\) to get this \[\Large \frac{7}{\sqrt{2}}\] \[\Large \frac{7\sqrt{2}}{\sqrt{2}*\sqrt{2}}\] \[\Large \frac{7\sqrt{2}}{\sqrt{2*2}}\] \[\Large \frac{7\sqrt{2}}{\sqrt{4}}\] \[\Large \frac{7\sqrt{2}}{2}\]